# Compress, Gather, and Recompute: REFORMing Long-Context Processing in Transformers

### v1

@@ Abstract
As large language models increasingly gain popularity in real-world applications, processing extremely long contexts, often exceeding the model’s pre-trained context limits, has emerged as a critical challenge. While existing approaches to efficient long-context processing show promise, recurrent compression-based methods struggle with information preservation, whereas random access approaches require substantial memory resources. We introduce REFORM, a novel inference framework that efficiently handles long contexts through a two-phase approach. First, it incrementally processes input chunks while maintaining a compressed KV cache, constructs cross-layer context embeddings, and utilizes early exit strategy for improved efficiency. Second, it identifies and gathers essential tokens via similarity matching and selectively recomputes the KV cache. Compared to baselines, REFORM achieves over 50% and 27% performance gains on RULER and BABILong respectively at 1M context length. It also outperforms baselines on $\infty$-Bench and MM-NIAH, demonstrating flexibility across diverse tasks and domains. Additionally, REFORM reduces inference time by 30% and peak memory usage by 5%, achieving both efficiency and superior performance.
Machine Learning
@@
1 Introduction
The ability to handle extremely long contexts, often exceeding the original model’s pre-trained context limits, has emerged as a critical challenge for the advanced usage of large language models (LLMs) in real-world scenarios. This capability is essential for various applications, such as processing life-long user interactions, understanding and debugging repository-level codebases, and handling multi-modal inputs (interleaved sequences of text and visual information can result in extremely long contexts). However, under existing Transformer-based language model architectures (Dubey et al., 2024; Jiang et al., 2023), processing such long sequences often causes significant computational challenges, requiring substantial computation as well as memory resources. These demanding requirements often prove infeasible in practical deployment settings, necessitating new technologies that can handle extremely long sequences with reasonable computational resources.
Current approaches to efficient infinite-context processing broadly fall into two categories: recurrent context processing and random access mechanisms. Recurrent methods (Xiao et al., 2023a; Zhang et al., 2023; Oren et al., 2024; Kim et al., 2024) divide the input into manageable chunks and iteratively process them while maintaining a summarized representation of prior chunks, typically by compressing or evicting parts of the Key-Value (KV) cache. While these approaches reduce memory and computational costs, they often suffer from ‘forgetting’ due to the loss of critical information during compression and/or eviction.
In contrast, another line of work aims to enable dynamic random-access to the previous inputs by preserving the full KV cache and retrieving relevant portions when processing new chunks (Xiao et al., 2024a; Liu et al., 2024). These methods provide more flexibility in accessing prior context, as they allow selective re-attention to specific segments of the input. However, maintaining the full KV cache requires substantial memory resources, often leading to significant memory overhead and latency increases, especially in practical deployments where CPU memory offloading is necessary. Furthermore, the increased flexibility does not necessarily lead to high retrieval performance. These limitations highlight the need for a more balanced approach that combines efficiency with precise long-context handling.
Figure 1: An overview of the proposed framework. REFORM efficiently processes long inputs through two phases. In the recurrent chunked forwarding phase, it segments inputs into chunks and processes them iteratively. In each iteration, REFORM (1) forwards each chunk conditioned on the previous KV cache, (2) extracts key QKV states from selected layers and heads for constructing cross-layer context embeddings, and (3) compresses the cache via token eviction (Zhang et al., 2023). An early exit strategy skips upper layers beyond those used for embedding collection, further improving efficiency. In the on-demand cache recomputation phase, REFORM selects important tokens via cosine similarity search with the query embeddings (last part of the input), gathers them, and recomputes the KV cache for further generation.
To address the above challenges, we propose REFORM (REcurrent chunked Forwarding with On-demand cache RecoMputation), a novel inference framework that combines the efficiency of recurrent approaches with the superior recall capabilities of random-access methods. Our approach implements a computationally efficient compress-gather-recompute pipeline that processes long contexts through two key phases, as illustrated in Figure 1.
In the encoding phase, we process input tokens in chunks through an adaptive caching mechanism called recurrent chunked forwarding: as each chunk is processed, tokens are added to the KV cache and sparsified by retaining only the heavy hitters (most influential tokens). Using this progressively sparsified KV cache, we compute representations up to an intermediate transformer layer L, collecting QKV states from multiple layers and heads to generate and store lightweight cross-layer context embeddings for all tokens. This multi-faceted efficiency strategy—combining chunked processing, sparse KV cache updates, and early exit—significantly reduces both computation time and memory overhead, as we maintain only small representations for retrieval while dynamically managing KV cache sparsity.
In the recomputation phase, the query tokens (corresponding to the recent context) identify relevant historical tokens through cosine similarity matching with the stored retrieval embeddings, and only these selected tokens undergo full KV cache recomputation across all layers. While this phase requires full computation for selected tokens, this recomputation is crucial: it restores high-fidelity representations for contextually important tokens, ensuring accurate processing of long-range dependencies. By selectively recomputing only the most relevant tokens, we achieve an optimal balance between computational efficiency and model performance, allowing detailed historical context access while avoiding the costs of maintaining full representations for all tokens.
Our extensive evaluations demonstrate REFORM’s effectiveness across various long-context understanding tasks. In needle-in-a-haystack tests, REFORM achieves perfect recall for contexts up to 1 million tokens at various depths. On more complex benchmarks, REFORM significantly outperforms existing methods, achieving over 50% performance gain on RULER and 27% on BABILong at 1M context lengths with the Mistral-NeMo-Instruct-2407 (Jiang et al., 2023) model, compared to the best-performing baselines. On more realistic tasks from $\infty$-Bench, REFORM achieves 31.2% average performance, substantially exceeding the baseline performance of 24.8%.
Operating at the transformer architecture level, our method is modality-agnostic and applicable to any domain/modality the base model supports, making it particularly valuable for real-world applications requiring extensive context understanding. To demonstrate the flexibility of our method, we evaluate in on three multi-modal needle-in-a-haystack datasets, achieving 57.51% performance with the Pixtral-12B-2409 (Agrawal et al., 2024) model, exceeding the baseline performance of 52.95%.
REFORM also delivers substantial efficiency improvements over recent state-of-the-art long-context processing methods. Compared to InfLLM (Xiao et al., 2024a) and InfiniPot (Kim et al., 2024), REFORM reduces inference time by 80% and 33%, respectively, in evaluations with 256k token inputs. Additionally, REFORM operates with lower memory requirements, reducing peak memory usage by 32% and 5%, compared to InfLLM and InfiniPot, respectively. These results demonstrate that REFORM effectively combines the benefits of both recurrent compression and random access approaches while mitigating their respective limitations.
2 Related Works
In this section, we discuss the existing approaches to extend existing models to efficiently handle extremely long inputs. We categorize these methods into two groups: methods that use recurrent context processing and methods that allow random access. We provide a more comprehensive discussion on related works in Appendix A.
Recurrent context processing. To address the computational challenges of long-context processing, several studies explore the use of recurrence for greater efficiency. A line of works (Dai et al., 2019; Bulatov et al., 2022; Munkhdalai et al., 2024) introduce architectural changes to Transformers, enabling chunk-level recurrence operations to process long contexts in smaller, manageable units. However, these approaches typically necessitate extensive training of the model, and therefore is not directly applicable to existing pre-trained large language models. More recent efforts leverage KV cache eviction to iteratively encode input chunks and compress the KV cache, avoiding architectural modifications or additional training. For instance, StreamingLLM (Xiao et al., 2023a) maintains fluent generation by preserving initial and most recent tokens while compressing intermediate ones. Later approaches (Zhang et al., 2023; Oren et al., 2024; Kim et al., 2024) identify important tokens from the prior context, enabling more informative cache compression. Despite their efficiency, the process of compressing prior inputs often results in the loss of critical information, leading to ‘forgetting’ issues. Consequently, these methods may struggle with tasks requiring precise retrieval of earlier inputs.
Random access approaches. An alternative direction involves enabling random access to prior context, akin to full attention, but in a more computationally efficient manner. These methods typically store the full KV cache in memory and dynamically retrieve relevant tokens as needed. Some approaches train the model (Wu et al., 2022) or an auxiliary side-network (Wang et al., 2023b) to utilize the retrieved tokens effectively. More recently, training-free strategies have emerged, which store the full KV cache in memory and retrieve it dynamically (Xiao et al., 2024a; Liu et al., 2024). While these methods allow random access to any part of the input sequence, they introduce significant memory overhead due to the need to maintain large caches. In practice, this often necessitates CPU offloading, which can further increase latency. Furthermore, the flexibility to access previous context may not necessarily lead to high retrieval performance.
3 Method
Algorithm 1 Overview of REFORM
[…]
In this section, we present the details of our proposed method. In Section 3.1, we first describe REFORM’s recurrent chunk forwarding phase in detail. This phase efficiently constructs token-level, cross-layer context embeddings by segmenting the long input into multiple chunks and repeatedly processing them while conditioning on a compressed previous KV cache. In Section 3.2, we further elaborate on how we construct the cross-layer context embeddings. Finally, in Section 3.3, we describe how we use the context embeddings to identify the relevant input segments and highlight our on-demand cache recomputation framework that enables random access to previous contexts while maintaining the integrity of the KV cache. We outline the full procedure in Figure 1 and Algorithm 1.
3.1 Recurrent Chunked Forwarding with Early Exit
Recurrent chunked forwarding. To process extremely long inputs under limited computational budget and context windows, several KV cache compression methods (Xiao et al., 2023a; Zhang et al., 2023; Oren et al., 2024; Kim et al., 2024) suggest using an iterative approach. This approach repeatedly compresses the previous input’s KV cache and forwards a small portion of the remaining input based on the compressed cache. While many existing works demonstrate their context extrapolation capabilities by applying KV cache eviction when processing every token (Xiao et al., 2023a; Zhang et al., 2023), this process can become extremely slow for long inputs as it is not parallelizable. Instead, we segment the long input into larger chunks (32k tokens for our experiments) and apply KV cache compression after forwarding each chunk to better utilize parallel computation.
For KV cache compression, we employ attention-based token eviction following H2O (Zhang et al., 2023), which uses cumulative attention scores as a signal to determine the most significant context. We retain the tokens that receive the highest attention scores in the KV cache. After compression, we reassign the position IDs so that the tokens in the compressed cache have consecutive position IDs. This position reassignment allows the model to handle longer sequences beyond its pre-trained context limit.
The recurrent chunked forwarding approach enables the encoding of very long sequences with limited computational resources because it encodes all inputs conditioned on a limited KV cache. This bypasses the need to store a large KV cache corresponding to the long input or compute expensive self-attention on the long input. Meanwhile, it allows the model to encode the inputs in a way that’s aware of the previous context, thanks to the presence of such information in the compressed KV cache.
Early exit. Utilizing a compressive approach for creating embeddings introduces an additional benefit: efficiency can be further improved by employing an early exit strategy. As observed in Section 3.2, high-performing embeddings are often available in the lower Transformer layers. Therefore, forwarding the inputs through the remaining layers after the topmost layer used for embedding extraction is unnecessary and can be skipped. The proposed early exit strategy reduces both computation and memory requirements because we do not need to keep the KV cache for the upper layers.
3.2 Constructing Cross-Layer Context Embeddings
Prior research has revealed the existence of specialized Transformer heads distributed across different layers that can accurately retrieve relevant information from long context input (Wu et al., 2024). To construct informative embeddings, we thus analyze the retrieval performance of various heads and embeddings in Transformers to determine the most suitable ones for our method. Specifically, we compare the token-level retrieval performance of Transformer hidden states, attention scores (without positional encoding, to make it applicable to extremely long inputs), and the attention QKV states across Transformer layers.
Table 1: Comparing different LLM embeddings. Best-3 MNR scores (lower is better) corresponding to the hidden states and the attention states, measured by Mistral-Nemo-Instruct-2407. Scores are averaged over 500 synthetic multi-hop QA examples.
Type Dim. Top-1 Top-2 Top-3 Avg.
Hidden States 5120 9.40 9.63 9.80 9.61
Attention 160 6.91 7.70 7.81 7.47
Query 160 6.48 6.74 6.93 6.72
Key 160 6.77 7.31 7.41 7.16
Value 160 5.77 6.57 6.57 6.30
Figure 2: MNR Scores for Value Heads. The distribution of the MNR scores (lower is better) across value states of different attention heads, measured by Mistral-Nemo-Instruct-2407 model over 500 synthetic multi-hop QA examples. Recurrent chunked forwarding with 256-token heavy hitter budget was employed for computing the embeddings.
Embedding head identification. We conducted a set of experiments on a synthetically constructed multi-hop question-answering dataset, and used Mean Normalized Rank (MNR) scores to measure retrieval performance; see Section B.1 for details. We report the MNR scores for the top-performing layers and heads in Table 1. Interestingly, despite their smaller size, we observed that the retrieval performance of attention QKV states was often better than that of the commonly used LLM hidden states (Wang et al., 2023a). This finding suggests that, with careful selection of the appropriate heads, directly using the QKV states from the attention layer can achieve higher accuracy than the hidden states, while being more computationally efficient.
Figure 2 illustrates the distribution of the top-performing value heads. Notably, the best-performing heads are not necessarily from the final Transformer layers. This observation implies that forwarding inputs through the upper layers may not be required for constructing effective retrieval embeddings, serving as a key motivation for our early exit strategy described in the previous section.
To create a universal embedding that is useful for a wide range of tasks, we do additional experiments to identify heads that can effectively represent complex input patterns. Specifically, we create a synthetic key-value retrieval task, which involves embedding multiple sentences of the format “The value corresponding to the id {key} is {value}.” within the WikiText (Merity et al., 2016) corpus, where keys and values are random 10-character ASCII strings.
We selected the top-performing embeddings for each synthetic dataset after evaluating 500 samples each. It is worth noting that although head selection is based on relatively short synthetic data (8k tokens), the benefits extrapolate to longer contexts involving millions of tokens.
$e_{\mathtt{comb}}=\mathtt{concat}\left(\left\{\frac{e_{\mathtt{i}}}{||e_{\mathtt{i}}||},\mathtt{i}\in\mathtt{selected\_heads}\right\}\right)$

### v2

@@ Abstract
As large language models increasingly gain popularity in real-world applications, processing extremely long contexts, often exceeding the model’s pre-trained context limits, has emerged as a critical challenge. While existing approaches to efficient long-context processing show promise, recurrent compression-based methods struggle with information preservation, whereas random access approaches require substantial memory resources. We introduce REFORM, a novel inference framework that efficiently handles long contexts through a two-phase approach. First, it incrementally processes input chunks while maintaining a compressed KV cache, constructs cross-layer context embeddings, and utilizes early exit strategy for improved efficiency. Second, it identifies and gathers essential tokens via similarity matching and selectively recomputes the KV cache. Compared to baselines, REFORM achieves over 52% and 34% performance gains on RULER and BABILong respectively at 1M context length. It also outperforms baselines on $\infty$-Bench, RepoEval, and MM-NIAH, demonstrating flexibility across diverse tasks and domains. Additionally, REFORM reduces inference time by 30% and peak memory usage by 5%, achieving both efficiency and superior performance.
00footnotetext: Work done during an internship at Amazon.
1 Introduction
The ability to handle extremely long contexts, often exceeding the original model’s pre-trained context limits, has emerged as a critical challenge for the advanced usage of large language models (LLMs) in real-world scenarios. This capability is essential for various applications, such as processing life-long user interactions, understanding and debugging repository-level codebases, and handling multi-modal inputs (interleaved sequences of text and visual information can result in extremely long contexts). However, under existing Transformer-based language model architectures dubey2024llama; jiang2023mistral, processing such long sequences often causes significant computational challenges, requiring substantial computation as well as memory resources. These demanding requirements often prove infeasible in practical deployment settings, necessitating new technologies that can handle extremely long sequences with reasonable computational resources.
Current approaches to efficient context window extrapolation broadly fall into two categories: recurrent context processing and random access mechanisms. Recurrent methods (xiao2023efficient; zhang2023h2o; oren2024transformers; kim2024infinipot) divide the input into manageable chunks and iteratively process them while maintaining a summarized representation of prior chunks, typically by compressing or evicting parts of the Key-Value (KV) cache. While these approaches reduce memory and computational costs, they often suffer from ‘forgetting’ due to the loss of critical information during compression and/or eviction.
Figure 1: An overview of the proposed framework. REFORM efficiently processes long inputs through two phases. In the recurrent chunked forwarding phase, it segments inputs into chunks and processes them iteratively. In each iteration, REFORM (1) forwards each chunk conditioned on the previous KV cache, (2) extracts key QKV states from selected layers and heads for constructing cross-layer context embeddings, and (3) compresses the cache via token eviction (zhang2023h2o). An early exit strategy skips upper layers beyond those used for embedding collection, further improving efficiency. In the on-demand cache recomputation phase, REFORM selects important tokens via similarity search with the query embeddings (last part of the input), gathers them, and recomputes the KV cache for further generation.
In contrast, another line of work aims to enable dynamic random-access to the previous inputs by preserving the full KV cache and retrieving relevant portions when processing new chunks (xiao2024infllm; liu2024reattention). These methods provide more flexibility in accessing prior context, as they allow selective re-attention to specific segments of the input. However, maintaining the full KV cache requires substantial memory resources, often leading to significant memory overhead and latency increases, especially in practical deployments where CPU memory offloading is necessary. Furthermore, the increased flexibility does not necessarily lead to high retrieval performance. These limitations highlight the need for a more balanced approach that combines efficiency with precise long-context handling.
To address the above challenges, we propose REFORM (REcurrent chunked Forwarding with On-demand cache RecoMputation), a novel inference framework that combines the efficiency of recurrent approaches with the superior recall capabilities of random-access methods via a computationally efficient compress-gather-recompute pipeline. In contrast to existing recurrent methods that use compressed KV cache for generation, REFORM uses compression to construct and store lightweight token-wise embeddings of the input. Given a query, REFORM then uses these embeddings to gather most relevant input tokens via similarity matching, and recomputes the full KV cache for those tokens. This process yields high-fidelity yet efficient representations for query-relevant tokens, leading to a superior retrieval ability of the method while still benefiting from reduced memory overhead.
Figure˜1 illustrates the overall approach. In the encoding phase, we process input tokens in chunks through an adaptive caching mechanism called recurrent chunked forwarding: as each chunk is processed, tokens are added to the KV cache and compressed by retaining only the heavy hitters (most influential tokens). Using this progressively sparsified KV cache, we compute representations up to an intermediate transformer layer L, collecting QKV states from multiple layers and heads to generate and store lightweight cross-layer context embeddings for all tokens. This multi-faceted efficiency strategy—combining chunked processing, sparse KV cache updates, and early exit—significantly reduces both computation time and memory overhead, as we maintain only small representations for retrieval while dynamically managing KV cache sparsity.
In the recomputation phase, the query tokens (corresponding to the recent context) identify relevant historical tokens through similarity matching with the stored retrieval embeddings, and only these selected tokens undergo full KV cache recomputation across all layers. While this phase requires full computation for selected tokens, this recomputation is crucial: it restores high-fidelity representations for contextually important tokens, ensuring accurate processing of long-range dependencies. By selectively recomputing only the most relevant tokens, we achieve a better balance between computational efficiency and model performance, allowing detailed historical context access while avoiding the costs of maintaining full representations for all tokens.
Our extensive evaluations demonstrate REFORM’s effectiveness across various long-context understanding tasks. In needle-in-a-haystack tests, REFORM achieves perfect recall for contexts up to 1 million tokens at various depths. On more complex benchmarks, REFORM significantly outperforms existing methods, achieving over 52% performance gain on RULER and 34% on BABILong at 1M context lengths with the Mistral-NeMo-Instruct-2407 (jiang2023mistral) model, compared to the best-performing baselines. On $\infty$-Bench, REFORM achieves 50.2% average accuracy with the same model, substantially exceeding the baseline performance of 37.6%. REFORM also outperforms the baselines in RepoEval, scoring 65.3% performance with Qwen2.5-Coder-1.5B-Instruct on API-level code completion while the best baseline gives 61.8%.
Operating at the transformer architecture level, REFORM is modality-agnostic and applicable to any domain/modality the base model supports. To demonstrate its flexibility, we evaluate REFORM on three multi-modal needle-in-a-haystack datasets, achieving 57.51% performance with the Pixtral-12B-2409 (agrawal2024pixtral) model, exceeding the baseline performance of 52.95%.
Finally, REFORM delivers substantial efficiency improvements over recent state-of-the-art long-context processing methods. Compared to InfLLM xiao2024infllm and InfiniPot (kim2024infinipot), REFORM reduces inference time by 80% and 33% and peak memory usage by 32% and 5% respectively, in evaluations with 256k token inputs. These results demonstrate that REFORM effectively combines the benefits of both recurrent compression and random access approaches while mitigating their respective limitations.
2 Related Works
In this section, we discuss the existing approaches for extending LLM’s native context window to efficiently handle extremely long inputs. We categorize these approaches into two groups: methods that use recurrent context processing and methods that leverage random access. We provide a more comprehensive discussion on related works in Appendix˜B.
Recurrent context processing. To address the computational challenges of long-context processing, several studies explore the use of recurrence for greater efficiency. A line of works (dai2019transformer; bulatov2022recurrent; munkhdalai2024leave) introduce architectural changes to Transformers, enabling chunk-level recurrence operations to process long contexts in smaller, manageable units. However, these approaches typically necessitate extensive training of the model, and therefore is not directly applicable to existing pre-trained large language models. More recent efforts leverage KV cache eviction to iteratively encode input chunks and compress the KV cache, avoiding architectural modifications or altering the model parameters. For instance, StreamingLLM (xiao2023efficient) maintains fluent generation by preserving initial and most recent tokens while compressing intermediate ones. Later approaches (zhang2023h2o; oren2024transformers; kim2024infinipot) identify important tokens from the prior context, enabling more informative cache compression. Despite their efficiency, the process of compressing prior inputs often results in the loss of critical information, leading to ‘forgetting’ issues. Consequently, these methods may struggle with tasks requiring precise retrieval of earlier inputs. REFORM addresses this issue through its gather and recompute phases, which yields high-fidelity representation of all query-relevant input tokens.
Random access approaches. An alternative direction is to enable random access to prior context, akin to full attention, but in a more computationally efficient manner. These methods typically store the full KV cache in memory and dynamically retrieve relevant tokens as needed. Some approaches train the model (wu2022memorizing) or an auxiliary side-network (wang2023augmenting) to utilize the retrieved tokens effectively. More recently, several strategies achieve the same goal without modifying the model parameters, by storing the full KV cache in memory and retrieving it dynamically (xiao2024infllm; liu2024reattention). While these methods allow random access to any part of the input sequence, they introduce significant memory overhead due to the need to maintain large caches. In practice, this often necessitates CPU offloading, which can further increase latency. Furthermore, the flexibility to access previous context may not necessarily lead to high retrieval performance. In contrast, REFORM uses KV cache compression and constructs compact token-level embeddings using only the high-performing heads to reduce memory overhead while still maintaining high retrieval performance.
3 Method
In this section, we present the details of our proposed method. In Section˜3.1, we first describe REFORM’s recurrent chunk forwarding phase in detail. This phase efficiently constructs token-level, cross-layer context embeddings by segmenting the long input into multiple chunks and repeatedly processing them while conditioning on a compressed previous KV cache. In Section˜3.2, we further elaborate on how we construct the cross-layer context embeddings. Finally, in Section˜3.3, we describe how we use the context embeddings to identify the relevant input segments and highlight our on-demand cache recomputation framework that enables random access to previous contexts while maintaining the integrity of the KV cache. We outline the full procedure in Figure˜1 and Algorithm˜1.
3.1 Embedding Extraction with Recurrent Chunked Forwarding and Early-Exit
Embedding extraction with recurrent chunked forwarding. To process extremely long inputs under limited computational budget and context windows, we adopt an iterative KV cache compression approach (xiao2023efficient; zhang2023h2o; oren2024transformers; kim2024infinipot). Specifically, we segment the long input into larger chunks (32k tokens for our experiments) and apply KV cache compression after forwarding each chunk to better utilize parallel computation. For KV cache compression, we employ attention-based token eviction following H2O (zhang2023h2o). After compression, we reassign the position IDs so that the tokens in the compressed cache have consecutive position IDs. This position reassignment allows the model to handle longer sequences beyond its pre-trained context limit.
Unlike existing approaches that aim to directly use these compressed KV cache for generation, we use it only to construct context embeddings that will later be used to identify which part of the input is required for generating a response.
Early exit. Utilizing a compressive approach for creating embeddings introduces an additional benefit: efficiency can be further improved by employing an early exit strategy. As observed in Section˜3.2, high-performing embeddings are often available in the lower Transformer layers. Therefore, forwarding the inputs through the remaining layers after the topmost layer used for embedding extraction is unnecessary. The proposed early exit strategy reduces both computation and memory requirements because we do not need to keep the KV cache for the upper layers.
3.2 Constructing Cross-Layer Context Embeddings
Table 1: Comparing similarity search methods. Best-3 MNR scores (lower is better) corresponding to different similarity search methods, including attention, and cosine similarity search using hidden states (HS) or attention QKV states. Scores are measured with Mistral-Nemo-Instruct-2407, and averaged over 500 multi-hop QA examples.
Type Dim. Top-1 Top-2 Top-3 Avg. Attention 160 6.91 7.70 7.81 7.47 Cosine-HS 5120 9.40 9.63 9.80 9.61 Cosine-Q 160 6.48 6.74 6.93 6.72 Cosine-K 160 6.77 7.31 7.41 7.16 Cosine-V 160 5.77 6.57 6.57 6.30
Prior research has revealed the existence of specialized Transformer heads distributed across different layers that can accurately retrieve relevant information from long context input (wu2024retrievalhead). To construct informative embeddings, we thus analyze the retrieval performance of various heads and embeddings in Transformers to determine the most suitable ones for our method. Specifically, we compare the token-level retrieval performance of attention scores (without positional encoding, to make it applicable to extremely long inputs), and cosine similarity between hidden states, or the attention QKV states (i.e. embeddings resulting from QKV projection in attention layers) across Transformer layers.
Embedding head identification. We conducted a set of experiments on a multi-hop question-answering dataset to identify the best embeddings to use for the similarity search; see Section˜C.1 for details. We report the MNR scores for the top-performing layers and heads in Table˜1. Somewhat surprisingly, we observe that cosine similarity search with attention QKV states often outperform the widely-used attention scores. It also outperformed using cosine similarity between the hidden states despite having a much smaller size. This finding suggests that, with careful selection of the appropriate heads, directly using the QKV states from the attention layer and using them for cosine similarity search can achieve a very high performance, while requiring minimal memory.
Figure 2: MNR Scores for Value Heads. The distribution of the MNR scores (lower is better) across value states of different attention heads, measured by Mistral-Nemo-Instruct-2407 model for 500 synthetic multi-hop QA examples. 256-token heavy hitter budget was used for computation.
Figure˜2 illustrates the distribution of the top-performing value heads. Notably, the best-performing heads are not necessarily from the final Transformer layers. This observation implies that forwarding inputs through the upper layers may not be required for constructing effective retrieval embeddings, serving as a key motivation for our early exit strategy described in the previous section.
To create a universal embedding that is useful for a wide range of tasks, we do additional experiments to identify heads that can effectively represent complex input patterns. Specifically, we create a synthetic key-value retrieval task, which involves embedding multiple sentences of the format “The value corresponding to the id {key} is {value}.” within the WikiText (merity2016pointer) corpus, where keys and values are random 10-character ASCII strings.
We selected the top-performing embeddings for each synthetic dataset after evaluating 500 samples each. We highlight that although head selection is based on relatively short synthetic data (8k tokens), the benefits extrapolate to longer contexts involving millions of tokens.